• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name markowitz

    Log entry time 23:57:31 on April30,2003

    Entry number 99243

    keyword=shift summary

    NOTE: AT END OF SHIFT WE REALIZED WE WERE LOOKING AT NEGATIVE HELICITY ONLY. ALL SCALER RATES IN THIS SUMMARY NEED TO BE DOUBLED.

    17:50 MCC delivers tune beam to hall, 1-3 uA
    start harp scans. Noted OTR data invalid.
    17:58 beam trip
    18:35 beam back at 4 uA, ask MCC to optimize OTR.
    18:55 OTR measurement l0oks good.

    Run 1412, septum off with no target, rates with 4 uA are:
    40 Hz T1
    3 Hz T2
    178 Hz T3
    14 Hz T4

    19:10 inserting long foil.
    MCC goes to tune mode for hall C, no beam.
    17:40
    Run 1413, septum off, long foil inserted, 4 uA
    40 Hz T1
    5 Hz T2
    200 Hz T3
    10 Hz T4

    19:45 Phone MCC to remove target
    Preparing to ramp septum. Ion chambers read 0. LeRose is
    making a strip chart.
    19:55 start ramping septum
    run 1414 has the data letting us get scaler rates vs. (via epics) septum current. Empty target is in. At 60 A, T1 is 250 Hz other trigger rates unchanged.
    Magnet temperature (upper coil) starting to drop at 70 A, from 7K to 6.85 K. LeRose claims that to be signficant. Beam trip.

    20:15 Beam back. No target meaning He4 gas (not an empty reference cell.)
    T1 1200 Hz
    T2 20 Hz
    T3 200 Hz
    T4 10 Hz with 100 A in septum. Beam current is 4 uA

    20:25 MCC calls. A com-fault was traced to a bad crate which is being replaced. Beam gone for another 10 minutes.
    20:50
    T1 5000 Hz
    T2 40 Hz
    T3 208 Hz
    T4 11.9 Hz
    at 150 A in the septum.

    21:10 Septum at 200 Amps.
    T1 10,000 Hz
    T2 70 Hz
    T3 212 Hz
    T4 25 Hz

    21:26 Septum @ 250
    T1 10,000 Hz
    T2 50
    T3 200
    T4 25

    21:45 Septum at 300 A
    T1 8000
    T2 75
    T3 250
    T4 50 Hz. Still 4 uA of beam.
    Starting to see a rise in the beam dump ioniization chambers. See the separate beam stripline by John LeRose.

    22:00 MCC calls. They had to secure beam to do injector tests. Beam gone for a few minutes.

    22:30 Called MCC. It will be at least another 20 minutes. Meantime the septum is at 350 A.

    22:45 beam back. Final excitation: 350 A in septum, no target in path (just 4He gas).
    T1 4800
    T2 50
    T3 250
    T4 50 Hz. end run 1419 (long single run).

    Dropping septum to 225 A.
    22:58 Septum goes into uncontrolled downward current run to zero. However it did not quench and John LeRose thinks the recover may be short. Coil temperature (upper coil) stayed in the cold range.

    23:45 Septum is back to 225 A. We have 4 uA and 10 KHz of T1, so we drop to 1 uA of current and take empty target (no reference cell) data at 2.5 KHz of T1. Run is 1422.

    NOTE: AT END OF SHIFT WE REALIZE WE WERE ONLY LOOKING AT NEGATIVE HELICITY SCALERS ALL NIGHT. ALL SCALER RATES ON THIS REPORT SHOULD BE DOUBLED.